OpenStudy (anonymous):

What are the recursive and explicit formulas for the following sequence? Year 0: 10 Year 1: 30 Year 2: 90 Year 3: 270 Year 4: 810 Year 5: 2430

OpenStudy (campbell_st):

well the 1st term is 10 and then you multiply each term by 3 this is a geometric series a term can be found by using \[a_{n} = a \times r^{n -1}\] you know a the 1st term is 10 and the know the terms are multiplied by 3 so you can substituet these values to get the formula where n >= 1 the alternative formula is \[a _{n + 1} = 3 \times a_{n} \] again n > = 1

OpenStudy (anonymous):

Thanks :D